# Cluster analysis of earthquake hypocenters in Azerbaijan and surrounding territories

### v1

@@ 2 The earthquake data selection
We examine earthquakes occurring within the rectangular region bounded by latitudes 38 to 42 N and longitudes 45 to 50 E. From the catalog [32], a total of 6,194 earthquakes with magnitudes $M\geq 2$ and depths exceeding 0.1 km were identified. A comparable dataset of Azerbaijani earthquakes has also been analyzed by [9]. Each seismic event is defined by its latitude, longitude, depth, time, and various magnitude types.
As a rule, information extracted from earthquake catalogs requires pre-processing [23], which typically involves the removal of non-tectonic events [24, 33] and those assigned characteristics by the catalog operator [33]. An essential step in catalog processing is the verification of its completeness, which includes the estimation of the completeness magnitude $M_{c}$ [9, 23].
This study does not involve filtering the catalog data. Instead, we directly proceed to estimate the completeness magnitude $M_{c}$, which represents the minimum magnitude above which all earthquakes are reliably recorded in the catalog. To determine $M_{c}$, we applied the MAXC method [34, 9]. Based on our previous analysis [8], a completeness magnitude of $M_{c}=2$ is appropriate, allowing us to include all earthquakes with $M\geq M_{c}$ in the analysis.
@@ 3 Morisita index application
Since our focus is on identifying earthquake groupings, it is appropriate to begin with a coarse criterion for clustering. To this end, we apply an approach that estimates the dispersion of spatially distributed points. Specifically, we use the box-counting method, in which the study area is divided into small cells—typically square grids—and the number of points within each cell is counted. This provides a basis for introducing the Morisita Index [35, 36, 37], a quantitative measure of spatial clustering.
@@
where $Q$ is the number of grid cells (boxes) covering the study area, $n_{i}$ denotes the number of points within the $i$th cell, and $N$ is the total number of points in the area.
The Morisita Index can be formally extended to multi-dimensional data. However, the computational complexity increases significantly with dimensionality. Therefore, it is practical to perform preliminary calculations using two-dimensional data representing earthquake epicenter locations, and subsequently compare the results with those obtained from three-dimensional data (hypocenter coordinates), where applicable.
@@
where $\{P_{\min},P_{\max}\}=\{41.700;52.128\}$, $\{R_{\min},R_{\max}\}=\{37.177;42.482\}$, $\{H_{\min}$, $H_{\max}\}=\{0;112.1\}$. Then each side of the cube is divided into $n$ parts of the length $\delta=1/n$.
@@
We begin the MI calculation with the case of 2D data, which, after the transformation (2), is distributed over the unit square, as shown in Fig.1b. The unit square is subdivided into $Q=n^{2}$ quadrats. The resulting Morisita index plot is presented in Fig. 2a (the lower broken line), illustrating the trend as the number $n$ increases from 1 to 100, or equivalently, as the mesh size $\delta$ decreases from 1 to 1/100.
The graph is bounded on the right by the value $I_{\delta}=1$ at $\delta=1$ (or $n=1$). In particular, local minima (”pits”) are observed at $n=m_{1}=8$, $n=m_{2}=12$, etc. To investigate the nature of these features, auxiliary studies were conducted.
To better understand the emergence of ”pits”, we compare the positions of two meshes obtained at $n=m_{2}=12$ and $n=11$, demonstrating their relation to cluster capture. The zoomed-in section of the unit square from Fig. 1b, which contains both meshes and a cluster, is shown in Fig. 2b.
Consider the partial quadrat formed by the less dense mesh, assuming that it contains $r$ cluster points, with points outside the square excluded from consideration. The denser mesh subdivides this quadrat into four smaller quadrats, each containing fewer cluster points—denoted as $r_{1,2,3,4}$—such that $r=\sum r_{i}$. Then the expression (1) for $I_{\delta}^{11}$ incorporates the term $r(r-1)$, while $I_{\delta}^{12}$ contains $\sum r_{i}(r_{i}-1)$. Let us examine their difference $r(r-1)-\sum r_{i}(r_{i}-1)=r^{2}-\sum r_{i}^{2}>0$. From this it follows that $I_{\delta}^{11}>I_{\delta}^{12}$. The above considerations offer a qualitative framework for understanding how the Morisita Index effectively distinguishes a sufficiently contrasting cluster from a background of sparsely distributed points.
(a)                                                                       (b)
Figure 2: a: Morisita Index $I_{\delta}$ vs. $\delta$. Here $m_{1}=8$ and $m_{2}=12$. Inset shows the sequence $L_{n}=\ln I_{1/n}/\ln n$, defined in (3), approaching its constant value (green). The plots are evaluated for 2D (black) and 3D (red) data, respectively. b: The part of Fig.1b with meshes corresponding to $\delta=1/m_{2}=1/12$ (red) and $\delta=1/11$ (green).
Transforming the locations of earthquake hypocenters using relation (2), we perform similar MI calculations for the 3D data, which allows us to construct the upper broken line in Fig. 2a. Unfortunately, the calculation of $I_{\delta}$ was terminated after $n=30$ due to the high computational demands. Nevertheless, comparing the shapes of the $I_{\delta}$ curves obtained for the 3D and 2D data reveals a clear similarity. Moreover, the ”pits” become more pronounced, although their positions remain unchanged.ronounced, while their positions remain unchanged.
@@
Thus, since the MI exceeds 1, we can conclude that earthquakes are distributed non-uniformly and tend to form clusters. Before proceeding to cluster identification, let us consider another feature of MI for the sake of completeness.
@@
Moreover, MI analysis provides additional insight into the fractal nature of datasets [23]. The method of MI calculation—namely, the box-counting approach—is also commonly used to assess the dimensions of fractal objects. Therefore, it is not surprising that MI is related to the fractal dimensions of self-similar point patterns. As shown by [36], the relationship between MI and the fractal dimension is established through an asymptotic power-law dependence, $I_{\delta}\propto\delta^{p}$ ($p<0$). According to [36], the exponent $p$ is related to the second-order Morisita slope (also referred to as the codimension of order 2) by $p=-S_{2}$. Furthermore, the relation $D_{2}-E=p$ holds, where $D_{2}$ is the second-order Rényi generalized dimension (or correlation dimension) and $E$ is the dimension of the Euclidean space in which the points are embedded. In addition, the following relation is valid (with $\delta=1/n$):
$\lim_{n\to\infty}L_{n}=\lim_{n\to\infty}\frac{\ln I_{1/n}}{\ln n}=p=const.$ (3)
To verify property (3), we evaluate the sequence $L_{n}$ for sufficiently large $n$, specifically for $n>10$. The resulting dependencies are shown in the inset of Fig. 2a. It can be observed that $L_{n}$ tends toward a constant value, with $p\approx-0.68$ for the 2D data and $p\approx-1.27$ for the 3D data. This indicates that the spatial distribution of earthquakes exhibits fractal properties.
@@
It is worth noting that [9] calculated the correlation dimension for spatially distributed epicenters to be approximately $D_{c}\approx 1.38$. Our evaluations yield a close estimate of the fractal dimension, with $D_{2}=2+p=1.32$. For the 3D case, the corresponding dimension is $D_{3}=3+p=1.73$.
@@ 4 DBSCAN and HDBSCAN application
To perform the cluster analysis, we use the well-known DBSCAN algorithm and its improved version, HDBSCAN, implemented via the sklearn.cluster module in Python [38]. Some preliminary results concerning the application of DBSCAN to this catalog were previously presented in [25]. Here, we summarize only the key points from [25].
To improve the results, we preprocess the data by scaling the earthquake depths according to the relation $Z=z/|z|$, where the Euclidean norm is defined as $|z|=\sqrt{\sum_{j}z_{j}^{2}}$.
@@
Next, the DBSCAN algorithm is applied to the catalog. Recall that DBSCAN ([26, 16, 22]) requires two intrinsic parameters: $\varepsilon$ and min_samples. The former defines the $\varepsilon$-neighborhood of each point, specifying the maximum distance between any two points within the neighborhood. The latter sets the minimum number of points required within the $\varepsilon$-neighborhood to form a cluster. To estimate appropriate ranges for these parameters, we apply the approach proposed by [26], with an adaptation detailed in Appendix A.
@@
Thus, the parameter $\varepsilon$ is selected from the range $\varepsilon\in[0.2,0.6]$ with a step size of 0.05, and min_samples is chosen from the range $[80,160]$ with a step size of 5. Two types of distance metrics are used—”euclidean” and ”manhattan”—resulting in a total of 306 cases.
@@
The result of DBSCAN is a partition consisting of clusters along with a set of unclassified points (noise points). It is important to note that different combinations of the parameters $\varepsilon$ and min_samples can yield partitions with the same number of clusters.s.
@@
The quality of point classification is assessed with the help of Silhouette Index [39], the set of which is shown in Fig.3a. To consider the influence of the parameter $\varepsilon$ to clustering, we associate point positions with a colorbar such that the lighter points correspond to higher values of $\varepsilon$. From Fig.3a it follows that increasing $\varepsilon$ leads to the reducing the number of clusters. A similar effect is observed with increasing min_samples.
@@
Number of clusters 4 5 6 7 8
$\varepsilon$ 0.5 0.4 0.45 0.45 0.4
min_samples 140 140 130 135 110
(a)                                                                       (b)
Figure 3: a: Silhouette Index vs. number of clusters. b: The 6-cluster partition corresponding to the maximal Silhouette Index depicted in the left panel by “$\times$”. The noise points are removed. Filled bullets mark earthquakes with $M>5.5$. The results obtained via DBSCAN.
(a)                                                                       (b)
Figure 4: a: Silhouette Index vs. number of clusters. b: The 6-cluster partition with noise points removed, corresponding to the maximal Silhouette Index, depicted in the left panel by the symbol “$\times$”. Filled bullets mark earthquakes with $M>5.5$. The results obtained via HDBSCAN.
Regarding the application of DBSCAN, it is worth making some remarks about the selection of the intrinsic parameter values. As shown in Fig.2a, MI profiles exhibit ”pits” that correspond to crucial changes in the structure of the dataset partition. In other words, MI can detect the approach of a cluster. This suggests that we can approximate the spatial scale for the mesh containing a cluster. For example, the first local pit is observed at $\delta=1/8$, and the quadrat containing the cluster has a side length of $\delta=1/7$. In the initial coordinates, the corresponding quadrat can have a side length of $\min{\delta(P_{\max}-P_{\min});\delta(R_{\max}-R_{\min})}=0.758$. The parameter $\varepsilon$ for DBSCAN defines a region with a side length of approximately $2\varepsilon$, which falls within the interval $(0.8;1.0)$ (Table 1) for cases where $4\leqslant N\leqslant 8$, and the SI approaches its maximum. There is also the case where $\varepsilon=0.3$ ([25]), which corresponds to a 7-cluster partition and lies close to SImax. Thus, the observed correlation between MI properties and DBSCAN parameters may not be accidental and can assist in applying the DBSCAN algorithm more effectively.
[…]
5 Adjusted Rand Index analysis
To quantify the differences between the partitions, we use the Adjusted Rand Index (ARI), which measures the similarity between two clusterings. Generally, these two partitions may consist of different numbers of subsets; however, in this study, we compare partitions containing the same number of clusters. It is important to note that the ARI ranges from $-1$ (indicating particularly discordant partitions) to 1 (indicating identical partitions). For the ARI evaluation, we use the adjusted_rand_score function from the sklearn.metrics module [40]. This approach helps us reduce the number of partitions that need to be considered as different.
In the following, we select partitions with a fixed number of clusters, $N$, which are compactly located and possess a sufficiently large SI. We begin with DBSCAN partitions and consider those with $N=5$ clusters, selecting the subset enclosed in the rectangle shown in Fig.3a. Similarly, partitions for $N=6$ and $N=7$ are also selected. The results of the ARI evaluation for pairs of clusterings obtained for $N=5$, $N=6$, and $N=7$ are presented in matrix form in Figs.5a, 5b, and 5c, respectively. The row and column labels correspond to the partitions for which the ARI is evaluated, listed in the order of their appearance within the specified rectangle in Fig.3a. The evaluated ARI values are placed in the corresponding cells and highlighted by color. ARIs for identical partitions are equal to 1 and are depicted in dark colors, while other combinations yield smaller ARIs, represented in lighter shades. We observe that the ARI varies within the range $[0.86;1]$ for the diagram in Fig.5a, within $[0.92;1]$ for the diagram in Fig.5b, and within $[0.72;1]$ for the diagram in Fig. 5c.
(a)                                                                       (b)
Figure 5: Adjusted Rand Index evaluation for 5- (a), 6- (b), and 7-cluster partition (c). The axis labels correspond to the selected partitions in Fig. 3a, numbered in descending order of SI, in particuler, the partition marked by “$\times$” correponds to the label 12 in the panel (b). The results obtained via DBSCAN.
(a)                                                                       (b)
Figure 6: Adjusted Rand Index evaluation for 5-cluster partition (a) and 6-cluster partition (b). The axis labels correspond to the selected partitions in Fig. 4a, numbered in descending order of SI, in particular, the partition marked by “$\times$” correponds to the label 0 in the panel (b). The results obtained via HDBSCAN.
Similarly, the ARI matrices are calculated for HDBSCAN partitions consisting of $N=5$ clusters and $N=6$ clusters, which are contained within the rectangles shown in Fig.4a. According to Fig.6a, the ARI fluctuates within the interval $[0.79;1]$ for the 5-cluster partitions. For the 6-cluster partitions, as shown in Fig.6b, the ARI takes values from the set ${0.86,0.87,0.88}$.
Further reduction in the number of partitions requires additional information about the earthquakes. In particular, we compare the clusters with the structure of the fault network. For the 6-cluster partitions obtained using both DBSCAN and HDBSCAN, we select the partitions that provide the maximum SI values. The locations of the clusters and the fault network are shown in Fig.3b and Fig.4b. A good agreement is observed between the cluster locations and the zones with denser and more diverse fault subsystems.
The selected partition should not be considered the only possible one. Other 6-cluster partitions, which we compared using ARI, can also be used.
6 Analysis of the earthquake depth distributions
@@
(e)                                                                       (f)
Figure 7: Histograms of the earthquake depth calculated for total number of earthquakes and earthquakes belonging to the specified cluster. Solid and dashed curves stand for the approximating PDFs. Clusters are taken from Fig.4b.

### v2

@@ 2 The earthquake data selection
We examine earthquakes occurring within the rectangular region bounded by latitudes 38 to 42 N and longitudes 45 to 50 E. From the catalog [32], a total of 6194 earthquakes with magnitudes $M\geq 2$ and depths exceeding 0.1 km were identified. A comparable dataset of Azerbaijani earthquakes has also been analyzed by [9]. Each seismic event is defined by its latitude, longitude, depth, time, and various magnitude types.
This study does not involve filtering the catalog data. We are interested in the spatial clustering of hypocenters and their depth distribution, which typically lie beyond the reach of regular human activity. Therefore, it is important to have access to the broadest possible set of earthquake hypocenters. Moreover, the catalog data for this region is quite uneven across years, sources, and methods of earthquake identification, and not sufficiently comprehensive. Thus, catalog pre-processing [23], which typically involves the removal of non-tectonic events [24, 33] and those assigned characteristics by the catalog operator [33], is not performed.
An essential step in catalog processing is the verification of its completeness, which includes the estimation of the completeness magnitude $M_{c}$ [9, 23], which represents the minimum magnitude above which all earthquakes are reliably recorded in the catalog. To determine $M_{c}$, we applied the MAXC method [34, 9]. Based on our previous analysis [8], a completeness magnitude of $M_{c}=2$ is appropriate, allowing us to include all 6194 earthquakes with $M\geq M_{c}$ in the analysis.
@@ 3 Morisita index application
Since our focus is on identifying earthquake groupings, it is appropriate to begin with a coarse criterion for clustering. To this end, we apply an approach that estimates the dispersion of spatially distributed points. Specifically, we use the box-counting method, in which the study area is divided into small cells, typically cubic grids, and the number of points within each cell is counted. This provides a basis for introducing a quantitative measure of spatial clustering in the form of the Morisita Index $I_{\delta}$, which, retaining the notation of works [35, 36, 37], is calculated as follows:
@@
where $\delta$ is the characteristic scale (side, diagonal, etc.) of a cell, $Q$ is the number of grid cells (boxes) covering the study area, $n_{i}$ denotes the number of points within the $i$th cell, and $N$ is the total number of points in the area. As mentioned in [35], $I_{\delta}=1$ when points are located completely randomly, $I_{\delta}<1$ is observed for uniformly or regularly dispersed points, and finally, $I_{\delta}>1$ means aggregated (clumped or clustered) point locations.
MI can be formally defined to multi-dimensional data. However, the computational complexity increases significantly with dimensionality. Therefore, it is practical to perform preliminary calculations using two-dimensional data representing earthquake epicenter locations, and subsequently compare the results with those obtained from three-dimensional data (hypocenter coordinates), where applicable.
@@
where $\{P_{\min},P_{\max}\}=\{42.3021,52.1283\}$, $\{R_{\min},R_{\max}\}=\{37.1765,43.2183\}$, $\{H_{\min}$, $H_{\max}\}=\{0;112.1\}$. Then each side of the cube is divided into $n$ parts of the length $\delta=1/n$.
@@
We begin the MI calculation with the case of 2D data, which, after the transformation (2), is distributed over the unit square, as shown in Fig.1b. The unit square is subdivided into $n^{2}$ quadrats. The resulting Morisita index plot is presented in Fig. 2 (the lower dashed curve), illustrating the trend as the mesh size $\delta$ decreases from 1 to 1/100. The graph is bounded on the right by the value $I_{\delta}=1$ at $\delta=1$.
Similarly, we perform MI calculations for the 3D data, which are substantially more computationally intensive, and plot the corresponding upper broken line in Fig. 2. Comparing the shapes of the $I_{\delta}$ curves obtained for the 3D and 2D data reveals a clear similarity.
@@
Thus, since the MI exceeds 1, we can conclude that earthquakes are not distributed entirely randomly and tend to form clusters. Before proceeding to cluster identification, let us consider several features of MI.
First, it is interesting to clarify the nature of “pits”, which are observed on the $I_{\delta}$ profiles (Fig.2), evaluated for both 2D and 3D data at $\delta=1/9$ (poorly distinguishable), $\delta=1/14$ (more visible), etc. Analyzing the mutual position of the mesh and the dataset, we noticed that for cell sizes $\delta=1/9$ or $\delta=1/14$ the mesh nodes fall inside some clusters, whereas for the larger values of $\delta$, i.e. 1/8 or 1/13, these clusters lie inside the mesh cell.
This can be explained qualitatively as follows. Consider the partial cell formed by the less dense mesh, assuming that it contains $r$ cluster points and neglecting the points outside this cell. The denser mesh subdivides this cell into four smaller cells, each containing fewer cluster points, denoted as $r_{1,2,3,4}$, such that $r=\sum r_{i}$. Then the expression (1) for $I_{\delta=1/8}$ incorporates the term r(r1)r(r-1), while $I_{\delta=1/9}$ contains $\sum r_{i}(r_{i}-1)$. Let us examine the difference between these terms and, by performing elementary transformations, prove that it is possitive, i.e., $r(r-1)-\sum r_{i}(r_{i}-1)=r^{2}-\sum r_{i}^{2}>0$. From this it follows that $I_{\delta=1/8}>I_{\delta=1/9}$. For 3D case, the mesh node falling into a cluster generates nine smaller mesh cubes. This reduces the number of cluster points that fall into smaller mesh cubes to a greater extent than in the 2D case. Therefore, the deeper and more pronounced “pits” are observed for the 3D case. The above considerations offer a qualitative framework for understanding how the Morisita Index effectively distinguishes a sufficiently contrasting cluster from a background of sparsely distributed points.
Figure 2: Morisita Index $I_{\delta}$ vs. $\delta$, represented in a logarithmic scale and evaluated for 2D (black dashed) and 3D (red solid) data, respectively.
@@
The second MI feature that we want to discuss concerns the asymptotic index behavior. MI analysis provides additional insight into the fractal nature of datasets [23]. The method of MI calculation, namely, the box-counting approach, is also commonly used to assess the dimensions of fractal objects. Therefore, it is not surprising that MI is related to the fractal dimensions of self-similar point patterns. As shown in [36], the relationship between MI and the fractal dimension is established through an asymptotic power-law dependence, $I_{\delta}\propto\delta^{p}$ ($p<0$). According to [36], the exponent $p$ is related to the second-order Morisita slope (also referred to as the codimension of order 2) by $p=-S_{2}$. Furthermore, the relation $D_{2}-E=p$ holds, where $D_{2}$ is the second-order Rényi generalized dimension (or correlation dimension) and $E$ is the dimension of the Euclidean space in which the points are embedded.
To verify this property, we consider the asymptotic behaviour of the sequence $I_{\delta}$ plotted in a logarithmic scale (Fig.2). To do this, the linear regression line y=0.1210.720xy=-0.121-0.720x evaluated by the least-squares method for the 2D data. Corresponding 95% confidence interval for the line slope is $\{-0.743,-0.697\}$ and the coefficient of determination is $R^{2}=0.975$. Thus, $p\approx-0.72$.
For the 3D data, the linear regression is y=0.1671.225xy=0.167-1.225x, 95% confidence interval for the slope $\{-1.28912,-1.1612\}$, and $R^{2}=0.982$. Thus, $p\approx-1.225$.
@@
It is worth noting that [9] calculated the correlation dimension for spatially distributed epicenters to be approximately $D_{c}\approx 1.38$. Our evaluations yield a close estimate of the fractal dimension, with $D_{2}=2+p=1.28$. For the 3D case, the corresponding dimension is $D_{3}=3+p=1.775$.
Thus, MI calculations show that the earthquake dataset exhibits fractal properties and is characterized by a pronounced cluster structure, for which a spatial scale can be assessed. Further studies concern the extraction of spatial clusters from the dataset utilizing the corresponding algorithms.
@@ 4 DBSCAN and HDBSCAN application
To perform the cluster analysis, we use the well-known DBSCAN algorithm and its improved version, HDBSCAN, implemented via the sklearn.cluster module in Python [38]. Some preliminary results concerning the application of DBSCAN to this catalog were previously presented in [25].
According to the recommendations of [16], the data prior to clustering should be scaled such that each component is transformed to a common range. We use transformation (2), which provides the translation of each coordinate into the range $(0;1)$.
@@
Next, the DBSCAN algorithm is applied to the scaled catalog. Recall that DBSCAN [26, 16, 22] requires two intrinsic parameters: $\varepsilon$ and min_samples. The former defines the $\varepsilon$-neighborhood of each point, specifying the maximum distance between any two points within the neighborhood. The latter sets the minimum number of points required within the $\varepsilon$-neighborhood to form a cluster. To estimate appropriate ranges for these parameters, we apply the approach proposed in [26], with an adaptation detailed in Appendix A. Note that both parameters influence the density and size of a cluster. Similarly, the Morishita index encapsulates information related to clustering and provides a rough estimate of cluster size. This parallel suggests a potential relationship between the DBSCAN parameters and the Morishita index. However, this subtle issue remains beyond the scope of these studies.
@@
Thus, the parameter $\varepsilon$ is selected from the range $\varepsilon\in[0.01,0.2]$ with a step size of 0.01, and min_samples is chosen from the range $[80,200]$ with a step size of 10. The euclidean distance metric is used, which for two points $x(x_{1},x_{2},x_{3})$ and $y(y_{1},y_{2},y_{3})$ is calculated as $d(x,y)=\sqrt{\sum_{i}^{3}{(x_{i}-y_{i})^{2}}}$.
@@
The result of DBSCAN is a partition consisting of clusters along with a set of unclassified points (noise points). It is important to note that different combinations of the parameters $\varepsilon$ and min_samples can yield partitions with the same number of clusters.
The quality of point classification is assessed with the help of Silhouette Index [39]. Let us recall that SI of the set $C$, which is partitioned by $N$ clusters $C_{i}$, $i=1,\dots,N$, is evaluated as follows $\text{SI}(C)=\sum_{i=1}^{N}s(i)/N$, where
$s(i)=\frac{a(i)-b(i)}{\max\{a(i),b(i)\}}.$
Similarly, the Silhouette Index is evaluated for each cluster $C_{i}$ according to the relation
$\text{SI}(C_{i})=\frac{1}{|C_{i}|}\sum_{j=1}^{|C_{i}|}s(j),$
where $|C_{i}|$ is the number of elements in the $C_{i}$ cluster. The mean intra-cluster distance $a(i)$ and mean nearest-cluster distance $b(i)$ are calculated via the following relations
$a(i)=\frac{1}{|C_{i}|-1}\sum_{j=1}^{|C_{i}|}d(i,j),\qquad b(i)=\min_{k\neq i}\sum_{j=1}^{|C_{k}|}d(i,j),$
where $d(i,j)$ is a distance.
When all clusters are well separated and dense enough, SI($C$) is close to 1. The low positive $\text{SI}(C)$ value indicates poorly separated clusters. When $1<\text{SI}(C)<0$, the elements refer to misclassified. In fact, SI technique provides good estimation of clustering quality when clusters are well localized and convex, and they should be linearly separable. SI doesn’t capture well the clusters of complex shapes or varying densities. In particular, such a situation occurs when DBSCAN algorithms are used. Then, to evaluate the clustering results, we employed various approaches and alternative indices [40, 41], including the evaluation of similarity between two partitions using the Adjusted Rand Index (ARI), which we apply below.
@@
Thus, the calculation of SI for each partition results in the diagram shown in Fig.3a. To consider the influence of the parameter $\varepsilon$ to clustering, we associate point positions with a colorbar such that the lighter points correspond to higher values of $\varepsilon$. From Fig. 3a it follows that SI decreases with increasing number of clusters in partitions. The parameters of partitions provided SImax are included in Table 1.
It is useful to find out the effect of both parameters $\varepsilon$ and min_samples on the number of clusters generated by DBSCAN. To do this, the diagram (Fig. 3b) representing the numbers of clusters in the partitions depending on $\varepsilon$ and min_samples is constructed on a dense grid, where $\varepsilon$ varies in the range $[0.01;0.15]$ with a step of 0.005 and min_samples ranges from $[10;600]$ with a step of 10. This shows that there is a bounded region in the two-parameter space whose points correspond to cases where there is more than one cluster in the partition. It has a complex structure, especially for small values of $\varepsilon$ and min_samples, and degenerates into a narrow band as the parameters increase.
Let us consider in more detail, the partitions with a medium number of clusters, e.g., cluster 4 (label “db 159”) and cluster 6 (label “db 192”) enclosed in rectangles in Fig.3a. The locations of the clusters are depicted in Figs.3c and d.
@@
Number of clusters 2 3 4 5 6 7
$\varepsilon$ 0.09 0.07 0.08 0.06 0.05 0.05
min_samples 190 120 200 150 80 90
@@
Figure 3: a: Silhouette Index vs. number of clusters. b: Two-parameter diagram showing the distribution of cluster counts. c and d: The 4- and 6-cluster partition marked in the Silhouette diagram by the labels “db 159” and “db 192”, respectively. The noise points are removed. Filled bullets mark strong earthquakes with $M>5.5$. The results obtained via DBSCAN.
(a)                   (b)
@@
(c)                   (d)
Figure 4: a: Silhouette Index vs. number of clusters. b, c, and d: The 4- and 6-cluster partitions with noise points removed, marked in the Silhouette diagram by labels “hdb 7”, “hdb 5”, and “hdb 15”, respectively. Filled bullets mark strong earthquakes with $M>5.5$. The results obtained via HDBSCAN.
Given that earthquake catalogs are often noisy and contain outliers, analyzing them with multiple algorithms is beneficial. For this reason, we also employ the HDBSCAN algorithm. It can work with clusters of varying densities, whereas DBSCAN tends to perform better with clusters of roughly uniform density. On the other hand, HDBSCAN may select clusters that are too dense and contain a small number of elements, that, in turn, reduces the possibility of conducting statistical analysis. Accordingly, we apply both methods and compare their clustering results.
Thus, let us perform similar evaluations of partitions using the HDBSCAN algorithm. We select the intrinsic parameter min_cluster_size from the range $(80,240)$ with a step of 5, while the parameter min_samples is set to its default value (equal to min_cluster_size), and either the euclidean or manhattan metric (i.e., $d(x(x_{1},x_{2},x_{3}),y(y_{1},y_{2},y_{3}))=\sum_{i}^{3}{|x_{i}-y_{i}|}$) is used. The SI values of each partition are calculated and the resulting points are shown in Fig.4a. In the partitions with the same number of clusters, SImax is achieved. For the specified parameter min_cluster_size range, the largest SImax is observed for 5-cluster partition, in contrast to the DBSCAN results (Fig.3a). For comparison with DBSCAN results, we consider the 4-cluster partition marked with the label of “hdb 7” (Fig.4a). The corresponding cluster location is plotted in Fig. 4b. We also consider two 6-cluster partitions labeled “hdb 5” and “hdb 15” in the diagram of Fig.4a, which possess nearly identical SIs but differ in the composition of several clusters (Fig. 4c and d).
5 Adjusted Rand Index analysis
To quantify the differences between the partitions, we use the Adjusted Rand Index (ARI) [41], which measures the similarity between two clusterings. Generally, these two partitions may consist of different numbers of subsets; however, in this study, we compare partitions containing the same number of clusters. It is important to note that the ARI ranges from 1-1 (indicating particularly discordant partitions) to 1 (indicating identical partitions). For the ARI evaluation, we use the adjusted_rand_score function from the sklearn.metrics module [42]. This approach helps us reduce the number of partitions that need to be considered as different.
In the following, we select partitions with a fixed number of clusters, $N$, which are compactly located in the rectangles depicted in the SI diagrams. We consider partitions with $N=4$ and $N=6$ clusters, selecting the subset enclosed in the rectangles shown in Fig.3a and Fig.4a. The results of the ARI evaluation for pairs of clusterings obtained for $N=4$ and $N=6$ are presented in matrix form in Figs.5a and 5b, respectively. The row and column labels correspond to the partitions, for which the ARI is evaluated, listed in the order of their appearance within the specified rectangles in Fig.3a and Fig.4a. Labels “db ” and “hdb” stand for partitions evaluated by DBScan and HDBScan method, respectively. The evaluated ARI values are placed in the corresponding cells and highlighted by color. ARIs for identical partitions are equal to 1 and are depicted in dark colors, while other combinations yield smaller ARIs, represented in lighter shades.
Analyzing Fig.5a, we see that 14 four-cluster partitions (labels “db 159 ”, $\dots$, “bd 212”) involve the set of similar partitions (labels “db 159” – “db 179”, “db 182” , “db 184”, and “db 189”) with ARI varying within the range $[0.72;1]$. However, there is another set of partitions, i.e. labels “db 181” and “db 185”, which are similar to each other with ARI=0.91, while similarity to other partitions varies in the range $[0.51;0.79]$. The group of partitions “db 203”– “db 212” consists of very similar partitions with ARI equal to 0.94 and 0.97, while they are similar to other partitions with the ARI range $[0.28;0.56]$ that means, in fact, weak similarity. Nine partitions produced by HDBSCAN (labels “hdb 7”, $\dots$, “hdb 23”) are also very similar to each other with ARI $\in[0.83;1]$. The comparison of pairs of partitions belonging to the sets of partitions obtained by the different algorithms shows that they differ significantly. For such cases, ARI varies from 0.49 to 0.15.
Interestingly, when we consider the 6-cluster partitions (Fig.5b), ARI indices are quite high both for the partitions obtained by the same algorithm and for the partitions produced by different ones. This can mean, in particular, that the algorithms we applied provide increasingly similar partitions as the number of clusters in them increases.
Figure 5: Adjusted Rand Index evaluation for 4-cluster (a) and 6-cluster partitions (b). The axis labels correspond to the selected partitions in Figs. 3a and 4a, numbered in descending order of SI. The labels with “db” and “hdb” stand for the results obtained via DBSCAN and HDBSCAN, respectively.
Further reduction in the number of partitions requires additional information about the earthquakes. In particular, we compare the clusters with the structure of the fault network. Moreover, we depicted earthquakes with magnitudes of M>>5.5 (filled bullets in Figs. 3c,d and 4b,c,d) representing the strongest earthquakes in this region and to some extent relate to local seismically active zones. For instance, as reported by [43] regarding the regions in Central Asia (Uzbekistan), the strong earthquakes (M $\geq$ 5) are often linked to the activation of deep faults within seismically active zones. Accordingly, we compare the localization of the obtained clusters with the spatial distribution of strong earthquakes.
For the 4-cluster partitions obtained using DBSCAN, the partition “db 159” can be selected as a representative of a set of similar 4-cluster partitions. This partition, depicted in Fig.3c, provides the maximum SI value (Fig.3a), its clusters are fairly equally filled with points, and cluster locations correlate well with fault consolidation and aggregation of strong earthquakes. Similar analysis allows one to select the partition “hdb 7” (Fig.4b). The ARI values for the pair “db 159” and “hdb 7” is 0.2 that tells us about the significant difference between partitions. Thus, depending on the aim of the investigation, we can select “db 159”, when the focus of further research is mainly on global issues, whereas “hdb 7” is more suitable for analyzing regions where earthquakes form dense, highly localized domains.
Considering in the similar manner 6-cluster partitions, we can select the partitions “db 192” and “hdb 15”. Each of them is a good representative of its class of similar partitions. They possess high values of SI and ARI of this pair is 0.48 that means a low degree of similarity. Analyzing the cluster locations of these partitions and the fault network (Fig.3d and Fig.4d), we see that cluster locations indeed differ, especially in the southern and southwestern parts of Azerbaijan. In particular, we use the partitions “db 192” and “hdb 15” for a comparative analysis of the depth distributions for Azerbaijani earthquakes.
6 Analysis of the earthquake depth distributions
(a) DBSCAN, “db 192”           (b) HDBSCAN, “hdb 15”
Figure 6: Fitted PDFs for depths corresponding to the clusters shown in a: Fig.3d and b: Fig.4d. The color of the curve corresponds to the color of the cluster depicted in the inset.